# A variational approach to dimension-free self-normalized concentration

### v1

@@ Appendix B Omitted Proofs
$\frac{D_{t}}{\lambda_{k(t)}}+\frac{\psi_{G,c}(\lambda_{k(t)})}{\lambda_{k(t)}}$ $=\frac{D_{t}}{\lambda_{k(t)}}+\frac{\lambda_{k(t)}}{2(1-c\lambda_{k(t)})}$
$\leq\frac{(1+c\sqrt{\frac{2D_{t}}{\alpha_{\delta,\eta}}})\sqrt{D_{t}}}{\sqrt{2/\alpha_{\delta,\eta}}}+\frac{\sqrt{2D_{t}}}{2(1+c\sqrt{2D_{t}})(1-c\frac{\sqrt{2D_{t}}}{1+c\sqrt{2D_{t}}})}$
$=\sqrt{\frac{\alpha_{\delta,\eta}D_{t}}{2}}+cD_{t}+\sqrt{\frac{D_{t}}{2}}$
$=cD_{t}+\left(\sqrt{\frac{\alpha_{\delta,\eta}}{2}}+\frac{1}{\sqrt{2}}\right)\sqrt{D_{t}}.$
Meanwhile, if $\lambda_{k(t)}=\psi_{G,c}^{-1}(1/\rho)$, then we may still use the lower bound in (B.20) and we obtain
$\frac{D_{t}}{\lambda_{k(t)}}+\frac{\psi_{G,c}(\lambda_{k(t)})}{\lambda_{k(t)}}$ $=\frac{D_{t}}{\lambda_{k(t)}}+\frac{1/\rho}{\psi_{G,c}^{-1}(1/\rho)}$
$\leq\frac{(1+c\sqrt{\frac{2D_{t}}{\alpha_{\delta,\eta}}})\sqrt{D_{t}}}{\sqrt{2/\alpha_{\delta,\eta}}}+\frac{c+\sqrt{c^{2}+2\rho}}{2\rho}$
$=\sqrt{\frac{\alpha_{\delta,\eta}D_{t}}{2}}+cD_{t}+\frac{c+\sqrt{c^{2}+2\rho}}{2\rho},$
@@
$\frac{D_{t}}{\lambda_{k(t)}}+\frac{\psi_{G,c}(\lambda_{k(t)})}{\lambda_{k(t)}}\leq\sqrt{\frac{\alpha_{\delta,\eta}D_{t}}{2}}+cD_{t}+\max\left\{\frac{c+\sqrt{c^{2}+2\rho}}{2\rho},\sqrt{D_{t}/2}\right\}.$ (B.21)
@@
$\psi^{-1}_{G,c}(\lambda_{k})\nearrow\frac{2}{c+\sqrt{c^{2}+2c}}.$ (B.22)
@@
$\alpha_{t}=\sup_{\theta\in\mathbb{R}^{d}}\frac{\langle\theta,U_{0}\theta\rangle}{\langle\theta,V_{t}\theta\rangle}\leq\frac{\sup_{\theta\in\mathbb{S}^{d-1}}\langle\theta,U_{0}\theta\rangle}{\inf_{\vartheta\in\mathbb{S}^{d-1}}\langle\vartheta,V_{t}\vartheta\rangle}=\frac{\gamma_{\max}(U_{0})}{\gamma_{\min}(V_{t})}.$
Therefore, for $t\in E_{k}$,
$g(\lambda_{k})$ $=\sqrt{\alpha_{t}\psi^{-1}_{G,c}(\lambda_{k})+1-\alpha_{t}}-\sqrt{\alpha_{t}\psi^{-1}_{G,c}(\lambda_{k})}$
$\geq 0\vee\left(\sqrt{1-\alpha_{t}}-\sqrt{\alpha_{t}\psi^{-1}_{G,c}(\lambda_{k})}\right)$
$\geq 0\vee\left(\sqrt{1-\frac{\gamma_{\max}(U_{0})}{\gamma_{\min}(V_{t})}}-\sqrt{\frac{\gamma_{\max}(U_{0})}{\gamma_{\min}(V_{t})}\frac{2}{(c+\sqrt{c^{2}+2c})}}\right)$
@@
$W_{c}(t,\lambda_{k(t)})=\frac{1}{\lambda_{k(t)}}\left(\frac{D_{t}}{\lambda_{k(t)}}+\frac{\psi_{G,c}(\lambda_{k(t)})}{\langle mbda_{k(t)}}\right)$ $\leq\frac{cD_{t}+\sqrt{\frac{\alpha_{\delta,\eta}}{2}D_{t}}+\max\left\{\frac{c+\sqrt{c^{2}+2\rho}}{2\rho},\sqrt{D_{t}/2}\right\}}{H_{c}(U_{0},V_{t})},$
@@
B.4 Proof of Proposition 4.7
@@
$\lambda^{*}(a)=\operatorname*{\textnormal{argmax}}_{\lambda\in[0,\lambda_{\max})}\lambda a-\psi(\lambda).$
@@
    For all $a\in\textnormal{Im}(\psi^{\prime})$, $a=\psi^{\prime}(\lambda^{*}(a))$ and $\lim_{a\to\infty}\lambda^{*}(a)=\lambda_{\max}$,
@@
We begin with (1). For fixed $a\in\textnormal{Im}(\psi^{\prime})$, let $f(x)=ax-\psi(x)$. Since $\psi$ is convex, $f$ is concave, so $f$ is maximized at $x^{*}$ satisfying $a=\psi^{\prime}(x^{*})$ ($x^{*}$ exists and is unique because $a\in\textnormal{Im}(\psi^{\prime})$ and $\psi^{\prime}$ is strictly increasing). But $x^{*}=\lambda^{*}(a)$ by definition, so $a=\psi^{\prime}(\lambda^{*}(a))$. Now, as $a\to\infty$, $\lambda=\lambda*(a)$ is chosen to maximize $a-\psi^{\prime}(\lambda)$, which grows as $\lambda\to\lambda_{\max}$ since $\psi^{\prime}$ is strictly increasing (since $\psi$ is strictly convex). (2) now also follows: since $\psi^{\prime}$ is strictly increasing it is invertible, so may invert (1) to get (2). For (3), note that $\psi^{\prime\prime\prime}\geq 0$ iff $\psi^{\prime\prime}$ is increasing iff $\psi^{\prime}$ is convex, hence its inverse $\lambda^{*}$ is concave (and nonnegative). Therefore, for all $x,y$, we have $\lambda^{*}(\alpha x+(1-\alpha)y)\geq\lambda^{*}\ell(x)+(1-\alpha)\lambda^{*}(y)$ for all $\alpha\in[0,1]$. Consider $y=0$, so $\lambda^{*}(\alpha x)\geq\alpha\lambda^{*}(x)+(1-\alpha)\lambda^{*}(0)\geq\alpha\lambda^{*}(x)$. Taking $x=ra$ and $\alpha=1/r\in[0,1]$, the result follows. For (4), if $h(\lambda)=\psi(\lambda)/\lambda^{2}$ is increasing, then $h^{\prime}(\lambda)\geq 0$, i.e.,
@@
Now let us return to the bound of Theorem 4.1. Let $\rho=\gamma_{\min}(U_{0})$. For all $\lambda\in[\psi^{-1}(1/\rho),\lambda_{\max})$, with probability $1-\delta$,
$\lambda\|S_{\tau}\|_{V_{\tau}^{-1}}-\psi(\lambda)\leq\frac{\|U_{0}\|_{{\textnormal{op}}}^{1/2}D_{\tau}(\delta)}{g_{\psi,\tau}(\lambda)}+\frac{g_{\psi,\tau}(\lambda)\psi(\lambda)}{\|U_{0}\|_{{\textnormal{op}}}^{1/2}}-\psi(\lambda).$
Since $g_{\psi,\tau}(\lambda)\leq 1$, if we assume that $U_{0}\succeq I_{d}$ (in which case $\|U_{0}\|_{{\textnormal{op}}}\geq 1$), then the above display implies that
$\lambda\|S_{\tau}\|_{V_{\tau}^{-1}}-\psi(\lambda)\leq\frac{\|U_{0}\|_{{\textnormal{op}}}^{1/2}D_{\tau}(\delta)}{g_{\psi,\tau}(\lambda)}.$ (B.23)
For $j\in\mathbb{N}$ and some $\eta>1$ to be determined, consider the event $E_{j}=\{\eta^{j}\leq\|S_{\tau}\|_{V_{\tau}^{-1}}<\eta^{j+1}\}$. Let $E_{0}=\{0\leq\|S_{\tau}\|_{V_{\tau}^{-1}}<2\}$. Note that $E_{0},E_{1},\dots$ partitions the sample space. Our strategy is to apply Theorem 4.1 once on each event $E_{j}$ with a different $\delta=\delta_{j}$ and $\lambda=\lambda_{j}$, where the latter can be optimized as (roughly) a function of $\|S_{\tau}\|_{V_{\tau}^{-1}}$ since we know how this quantity behaves on $E_{j}$. Rewriting (B.23) with $\delta=\delta_{j}$ and $\lambda=\lambda_{j}$ gives
$\lambda_{j}\|S_{\tau}\|_{V_{\tau}^{-1}}-\psi(\lambda_{j})\leq h_{\tau}(\lambda_{j})D_{\tau}(\delta_{j}),$ (B.24)
where $h_{\tau}(\lambda)=\|U_{0}\|_{{\textnormal{op}}}^{1/2}/g_{\psi,\tau}(\lambda)$. Now, conditioning on $E_{j}$, we have
$\lambda_{j}\eta^{j+1}-\psi(\lambda_{j})$ $=\lambda_{j}\|S_{\tau}\|_{V_{\tau}^{-1}}-\psi(\lambda_{j})+\lambda_{j}(\eta^{j+1}-\|S_{\tau}\|_{V_{\tau}^{-1}})$
$\leq h_{\tau}(\lambda_{j})D_{\tau}(\delta_{j})+\lambda_{j}(\eta^{j+1}-\eta^{j}).$ (B.25)
@@
$\lambda_{j}^{*}=\max\left\{\lambda^{*}(\eta^{j+1}),\psi^{-1}(1/\rho)\right\}.$ (B.26)
Let us suppose for now that $\lambda^{*}_{j}=\lambda^{*}(\eta^{j+1})$ so that $\lambda_{j}^{*}\eta^{j+1}-\psi(\lambda^{*}_{j})=\psi^{*}(\eta_{j+1})$ and we can rearrange (B.25) to read
$\psi^{*}(\|S_{\tau}\|_{V_{\tau}^{-1}})\leq\psi^{*}(\eta^{j+1})\leq h_{\tau}(\lambda_{j}^{*})D_{\tau}(\delta_{j})+\eta^{j}\lambda_{j}^{*}(\eta-1),$ (B.27)
where we’ve used that $\|S_{\tau}\|_{V_{\tau}^{-1}}\leq\eta^{j+1}$ on $E_{j}$ and the fact that $\psi^{*}$ is increasing. Now, we claim that
$\eta^{j}\lambda_{j}^{*}(\eta-1)\leq\frac{\psi^{*}(\eta^{j})}{2}.$
Let $s=\eta^{j}$ and define $R(s)=s\lambda_{j}^{*}(\eta-1)/\psi^{*}(s)$. We want to show that $R(s)\leq 1/2$. Using Lemma B.3, we have $\psi^{*}(s)\geq s\lambda^{*}(s)/2$. Moreover, if $\psi^{\prime\prime\prime}\geq 0$, then $\lambda^{*}(\eta s)\leq\eta\lambda^{*}(s)$, so
@@
$\psi^{*}(\|S_{\tau}\|_{V_{\tau}^{-1}})\leq h_{\tau}(\lambda_{j}^{*})D_{\tau}(\delta_{j})+\frac{\psi^{*}(\eta^{j})}{2}\leq h_{\tau}(\lambda_{j}^{*})D_{\tau}(\delta_{j})+\frac{\psi^{*}(\|S_{\tau}\|_{V_{\tau}^{-1}})}{2},$
whence $\psi^{*}(\|S_{\tau}\|_{V_{\tau}^{-1}})\leq 2h_{\tau}(\lambda_{j}^{*})D_{\tau}(\delta_{j})$, i.e.,
$\|S_{\tau}\|_{V_{\tau}^{-1}}\leq(\psi^{*})^{-1}\left(2h_{\tau}(\lambda_{j}^{*})D_{\tau}(\delta_{j})\right).$ (B.28)
Now we return to the scenario when $\lambda_{j}^{*}=\psi^{-1}(1/\rho)$ in (B.26). In this case, we have
$\|S_{\tau}\|_{V_{\tau}^{-1}}\leq\frac{h_{\tau}(\psi^{-1}(1/\rho))D_{\tau}(\delta_{j})+1/\rho}{\psi^{-1}(1/\rho)}.$
@@
$B_{\tau}=\begin{cases}(\psi^{*})^{-1}\left(2h_{\tau}(\lambda^{*}(\eta^{j(\tau)+1})D_{\tau}(\delta_{j(\tau)})\right),&\text{if }\lambda^{*}(\eta^{j(\tau)+1})\geq\psi^{-1}(1/\rho),\\ \frac{h_{\tau}(\psi^{-1}(1/\rho))D_{\tau}(\delta_{j})+1/\rho}{\psi^{-1}(1/\rho)},&\text{otherwise,}\end{cases}$ (B.29)
and $j(\tau)$ is the unique $j\in\mathbb{N}_{0}$ such that $\|S_{\tau}\|_{V_{\tau}^{-1}}\in[\eta^{j},\eta^{j+1})$. Therefore,
@@
Therefore, for large enough $t$, as long as $\psi^{-1}(1/\rho)<\lambda_{\max}$, we have
@@
$\|S_{\tau}\|_{V_{\tau}^{-1}}\leq B_{\tau}\lesssim(\psi^{*})^{-1}\left(2h_{\tau}(\lambda^{*}(\eta^{j(\tau)+1}))D_{\tau}(\delta_{j(\tau)})\right).$ (B.30)
It remains to bound $h_{\tau}(\lambda^{*}(\eta^{j(\tau)+1}))$. When $\lambda_{\max}<\infty$, we have that
$\lim_{j\to\infty}\psi^{-1}(\lambda^{*}(\eta^{j+1}))\nearrow\psi^{-1}(\lambda_{\max}).$
To see this, again start with the fact that $\lambda^{*}$ is increasing. Then, since $\psi$ is strictly increasing, $\psi^{-1}$ is continuous, so
$\lim_{j\to\infty}\psi^{-1}\left(\lambda^{*}(\eta^{j+1})\right)=\psi^{-1}\left(\lim_{j\to\infty}\eta^{j+1}\right)=\psi^{-1}(\lambda_{\max}).$
That the sequence increases to its limit follows since $\psi^{-1}$ is increasing. As was done in the proof of Proposition 4.3, notice that $\alpha_{t}\leq\gamma_{\max}(U_{0})/\gamma_{\min}(V_{t})$. Then,
$g_{\psi_{\tau}}(\lambda_{j}^{*})$ $=\sqrt{\alpha_{t}\psi^{-1}(\lambda^{*}(\eta^{j(\tau)+1}))+1-\alpha_{t}}-\sqrt{\alpha_{t}\psi^{-1}(\lambda^{*}(\eta^{j(\tau)+1}))}$
$\geq\sqrt{1-\alpha_{\tau}}-\sqrt{\alpha_{t}\psi^{-1}(\lambda_{\max})}$
$\geq\sqrt{1-\frac{\gamma_{\max}(U_{0})}{\gamma_{\min}(V_{\tau})}}-\sqrt{\frac{\gamma_{\max}(U_{0})\psi^{-1}(\lambda_{\max})}{\gamma_{\min}(V_{\tau})}}$
$\gtrsim\sqrt{1-\frac{1}{\gamma_{\min}(V_{\tau})}},$

### v2

@@ Appendix B Omitted Proofs
$\frac{D_{t}}{\lambda_{k(t)}}+\frac{\xi(c)\psi_{G,c}(\lambda_{k(t)})}{\lambda_{k(t)}}$ $=\frac{D_{t}}{\lambda_{k(t)}}+\frac{\xi(c)\lambda_{k(t)}}{2(1-c\lambda_{k(t)})}$
$\leq\frac{(1+c\sqrt{\frac{2D_{t}}{\alpha_{\delta,\eta}}})\sqrt{D_{t}}}{\sqrt{2/\alpha_{\delta,\eta}}}+\frac{\xi(c)\sqrt{2D_{t}}}{2(1+c\sqrt{2D_{t}})(1-c\frac{\sqrt{2D_{t}}}{1+c\sqrt{2D_{t}}})}$
$=\sqrt{\frac{\alpha_{\delta,\eta}D_{t}}{2}}+cD_{t}+\xi(c)\sqrt{\frac{D_{t}}{2}}$
$=cD_{t}+\left(\sqrt{\frac{\alpha_{\delta,\eta}}{2}}+\frac{\xi(c)}{\sqrt{2}}\right)\sqrt{D_{t}}.$
Meanwhile, if $\lambda_{k(t)}=\psi_{G,c}^{-1}(1/\upsilon)$, then we may still use the lower bound in (B.20) and we obtain
$\frac{D_{t}}{\lambda_{k(t)}}+\frac{\xi(c)\psi_{G,c}(\lambda_{k(t)})}{\lambda_{k(t)}}$ $=\frac{D_{t}}{\lambda_{k(t)}}+\frac{\xi(c)/\upsilon}{\psi_{G,c}^{-1}(1/\upsilon)}$
$\leq\frac{(1+c\sqrt{\frac{2D_{t}}{\alpha_{\delta,\eta}}})\sqrt{D_{t}}}{\sqrt{2/\alpha_{\delta,\eta}}}+\xi(c)\frac{c+\sqrt{c^{2}+2\upsilon}}{2\upsilon}$
$=\sqrt{\frac{\alpha_{\delta,\eta}D_{t}}{2}}+cD_{t}+\xi(c)\frac{c+\sqrt{c^{2}+2\upsilon}}{2\upsilon},$
@@
$\frac{D_{t}}{\lambda_{k(t)}}+\frac{\xi(c)\psi_{G,c}(\lambda_{k(t)})}{\lambda_{k(t)}}\leq\sqrt{\frac{\alpha_{\delta,\eta}D_{t}}{2}}+cD_{t}+\xi(c)\max\left\{\frac{c+\sqrt{c^{2}+2\upsilon}}{2\upsilon},\sqrt{\frac{D_{t}}{2}}\right\}.$ (B.21)
@@
$\psi^{-1}_{G,c}(\lambda_{k})\nearrow\frac{2}{c+\sqrt{c^{2}+2c}}=\xi(c).$ (B.22)
@@
$\alpha_{t}=\sup_{\theta\in\mathbb{R}^{d}}\frac{\langle\theta,U_{0}\theta\rangle}{\langle\theta,\widehat{V}_{t}\theta\rangle}\leq\frac{\sup_{\theta\in\mathbb{S}^{d-1}}\langle\theta,U_{0}\theta\rangle}{\inf_{\vartheta\in\mathbb{S}^{d-1}}\langle\vartheta,\widehat{V}_{t}\vartheta\rangle}=\frac{\gamma_{\max}(U_{0})}{\gamma_{\min}(\widehat{V}_{t})},$
and using the lower bound in (B.20),
$\psi^{-1}_{G,c}(\lambda_{k})$ $=\lambda_{k}(\sqrt{c^{2}+2c+2/\lambda_{k}}-c)$
$\geq\lambda_{k}(\sqrt{c^{2}+2c}-c)$
$\geq\frac{\sqrt{2D_{t}/\alpha_{\delta,\eta}}}{1+c\sqrt{2D_{t}/\alpha_{\delta,\eta}}}(\sqrt{c^{2}+2c}-c)$
$=\frac{\sqrt{c^{2}+2c}-c}{1/\sqrt{2D_{t}/\alpha_{\delta,\eta}}+c}$
$\geq\frac{\sqrt{c^{2}+2c}-c}{\sqrt{\alpha_{\delta,\eta}/\log(1/\delta)}+c}$
$\geq\frac{\sqrt{c^{2}+2c}-c}{\sqrt{\alpha_{\delta,\eta}}+c}.$
where we’ve used that $\log(1/\delta)\geq 1$ since $\delta\leq 1/e$. Therefore, for $t\in E_{k}$,
$g(\lambda_{k})$ $=\sqrt{\psi^{-1}_{G,c}(\lambda_{k})}-\sqrt{\alpha_{t}\psi^{-1}_{G,c}(\lambda_{k})}$
$\geq 0\vee\left(\sqrt{\psi^{-1}(\lambda_{k})}-\sqrt{\frac{\gamma_{\max}(U_{0})}{\gamma_{\min}(\widehat{V}_{t})}\xi(c)}\right)$
$\geq 0\vee\left(\sqrt{\frac{\sqrt{c^{2}+2c}-c}{\sqrt{\alpha_{\delta,\eta}}+c}}-\sqrt{\frac{\gamma_{\max}(U_{0})}{\gamma_{\min}(\widehat{V}_{t})}\xi(c)}\right)$
@@
$W_{c}(t,\lambda_{k(t)})$ $=\frac{1}{\lambda_{k(t)}}\left(\frac{D_{t}}{\lambda_{k(t)}}+\frac{\xi(c)\psi_{G,c}(\lambda_{k(t)})}{\lambda_{k(t)}}\right)$
$\leq\frac{cD_{t}+\sqrt{\frac{\alpha_{\delta,\eta}}{2}D_{t}}+\xi(c)\max\left\{\frac{c+\sqrt{c^{2}+2\upsilon}}{2\upsilon},\sqrt{D_{t}/2}\right\}}{H_{c}(U_{0},V_{t})},$
@@
B.4 Proof of Lemma 4.8
First consider $\psi=\psi_{G,c}$. From Lemma B.3 we know that $\psi^{\prime}(\lambda^{*}(a))=a$, i.e.,
$\frac{\lambda^{*}(a)(2-c\lambda^{*}(a))}{2(1-c\lambda^{*}(a))^{2}}=a.$
Solving for $\lambda^{*}(a)$ with the quadratic equation gives
$\lambda^{*}(a)=\frac{\sqrt{2ac+1}-1}{c\sqrt{2ac+1}},$
so
$\psi(\lambda^{*}(a_{n}))=\frac{(\sqrt{2ac+1}-1)^{2}}{2c^{2}\sqrt{2ac+1}}.$
As $a=a_{n}$ grows, we have $\psi(\lambda^{*}(a))=O(\sqrt{a_{n}}/c^{2})$. Meanwhile,
$\psi^{-1}(u)=\frac{2}{c+\sqrt{c^{2}+2/u}},$
so $\psi^{-1}(\lambda^{*}(a_{n}))\nearrow\psi^{-1}(1/c)$. Therefore,
$f(a_{n})\asymp\frac{\sqrt{a_{n}}}{c^{3/2}}\left(\sqrt{\frac{2}{c+\sqrt{c^{2}+2c}}}-1\right).$
Now, for $\psi=\psi_{E,c}$, setting $\psi^{\prime}(\lambda^{*}(a))=a$ gives
$\lambda^{*}(a)=\frac{a}{1+ca}.$
Hence
$\psi(\lambda^{*}(a))=\frac{\log(1+ca)}{c^{2}}-\frac{a}{c(1+ca)},$
and as $a=a_{n}$ grows with $n$, we have $\psi(\lambda^{*}(a_{n}))\lesssim\log(ca_{n}))$. Meanwhile, the inverse of $\psi$ obeys
$\psi^{-1}(u)=\frac{1+W_{0}(-e^{-1-uc^{2})}}{c},$ (B.23)
where $W_{0}$ is the principal branch of the W Lambert function, which is increasing, continuous, and satisfies $W_{0}(z)\exp(W_{0}(z))=z$ and has range $z\in[-1/e,\infty)$. Hence
$\psi^{-1}(\lambda^{*}(a))=\frac{1+W_{0}(-\exp(-1-\frac{c^{2}a}{1+a}))}{c}\nearrow\frac{1+W_{0}(-e^{-1-c})}{c}\leq 1/c.$
Therefore, $f_{\psi}(\lambda^{*}(a))\lesssim\log(ca_{n})(\sqrt{1/c}-1)$, which completes the argument.
B.5 Proof of Proposition 4.9
@@
$\lambda^{*}(a)\in\operatorname*{\textnormal{argmax}}_{\lambda\in[0,\lambda_{\max})}\lambda a-\psi(\lambda).$
@@
    $\lambda^{*}(a)$ is well-defined and unique. Further, for all $a\in\textnormal{Im}(\psi^{\prime})$, $a=\psi^{\prime}(\lambda^{*}(a))$ and $\lim_{a\to\infty}\lambda^{*}(a)=\lambda_{\max}$,
@@
We begin with (1). For fixed $a\in\textnormal{Im}(\psi^{\prime})$, let $f(x)=ax-\psi(x)$. Since $\psi$ is convex, $f$ is concave, so $f$ is maximized at $x^{*}$ satisfying $a=\psi^{\prime}(x^{*})$ ($x^{*}$ exists and is unique because $a\in\textnormal{Im}(\psi^{\prime})$ and $\psi^{\prime}$ is strictly increasing). But $x^{*}=\lambda^{*}(a)$ by definition, so $a=\psi^{\prime}(\lambda^{*}(a))$. Now, as $a\to\infty$, $\lambda=\lambda^{*}(a)$ is chosen to maximize $a-\psi^{\prime}(\lambda)$, which grows as $\lambda\to\lambda_{\max}$ since $\psi^{\prime}$ is strictly increasing (since $\psi$ is strictly convex). (2) now also follows: since $\psi^{\prime}$ is strictly increasing it is invertible, so may invert (1) to get (2). For (3), note that $\psi^{\prime\prime\prime}\geq 0$ iff $\psi^{\prime\prime}$ is increasing iff $\psi^{\prime}$ is convex, hence its inverse $\lambda^{*}$ is concave (and nonnegative). Therefore, for all $x,y$, we have $\lambda^{*}(\alpha x+(1-\alpha)y)\geq\lambda^{*}\ell(x)+(1-\alpha)\lambda^{*}(y)$ for all $\alpha\in[0,1]$. Consider $y=0$, so $\lambda^{*}(\alpha x)\geq\alpha\lambda^{*}(x)+(1-\alpha)\lambda^{*}(0)\geq\alpha\lambda^{*}(x)$. Taking $x=ra$ and $\alpha=1/r\in[0,1]$, the result follows. For (4), if $h(\lambda)=\psi(\lambda)/\lambda^{2}$ is increasing, then $h^{\prime}(\lambda)\geq 0$, i.e.,
@@
Now let us return to the bound of Theorem 4.1. Let $\widehat{V}_{t}=V_{t}+U_{0}$ and let $\upsilon=\gamma_{\min}(U_{0})$. For all $\lambda\in[\psi^{-1}(1/\upsilon),\lambda_{\max})$, with probability $1-\delta$, we can rearrange the guarantee to give
$\lambda\|S_{\tau}\|_{\widehat{V}_{\tau}^{-1}}-\psi(\lambda)\leq\frac{D_{\tau}(\delta)}{g_{\psi,\tau}(\lambda)}+g_{\psi,\tau}(\lambda)\psi(\lambda)-\psi(\lambda),$ (B.24)
where as usual
$D_{t}(\delta)=\frac{1}{2}\log\left(\frac{\det\widehat{V}_{t}}{\det U_{0}}\right)+1+\log(1/\delta).$
For $j\in\mathbb{N}$ and some $\eta>1$ to be determined, consider the event $E_{j}=\{\eta^{j}\leq\|S_{\tau}\|_{\widehat{V}_{\tau}^{-1}}<\eta^{j+1}\}$. Let $E_{0}=\{0\leq\|S_{\tau}\|_{\widehat{V}_{\tau}^{-1}}<\eta\}$. Note that $E_{0},E_{1},\dots$ partitions the sample space. Our strategy is to apply Theorem 4.1 once on each event $E_{j}$ with a different $\delta=\delta_{j}$ and $\lambda=\lambda_{j}$, where the latter can be optimized as (roughly) a function of $\|S_{\tau}\|_{\widehat{V}_{\tau}^{-1}}$ since we know how this quantity behaves on $E_{j}$. Rewriting (B.24) with $\delta=\delta_{j}$ and $\lambda=\lambda_{j}$ gives
$\lambda_{j}\|S_{\tau}\|_{\widehat{V}_{\tau}^{-1}}-\psi(\lambda_{j})\leq h_{\tau}(\lambda_{j})D_{\tau}(\delta_{j})+\psi(\lambda_{j})(g(\lambda_{j})-1),$ (B.25)
where $g(\lambda)=g_{\psi,\tau}(\lambda)$ and $h_{\tau}(\lambda)=1/g(\lambda)$. Conditioning on $E_{j}$, we have
$\lambda_{j}\eta^{j+1}-\psi(\lambda_{j})$ $=\lambda_{j}\|S_{\tau}\|_{\widehat{V}_{\tau}^{-1}}-\psi(\lambda_{j})+\lambda_{j}(\eta^{j+1}-\|S_{\tau}\|_{\widehat{V}_{\tau}^{-1}})$
$\leq h_{\tau}(\lambda_{j})D_{\tau}(\delta_{j})+\psi(\lambda_{j})(g(\lambda_{j})-1)+\lambda_{j}(\eta^{j+1}-\eta^{j}).$ (B.26)
@@
$\lambda_{j}=\max\left\{\lambda^{*}(\eta^{j+1}),\psi^{-1}(1/\upsilon)\right\}.$ (B.27)
Let us suppose for now that $\lambda_{j}=\lambda^{*}(\eta^{j+1})$ so that $\lambda_{j}\eta^{j+1}-\psi(\lambda_{j})=\psi^{*}(\eta^{j+1})$ (by definition of $\psi^{*}$). In this case we have
$\psi(\lambda_{j})(g(\lambda_{j})-1)\leq\psi(\lambda_{j})\left(\sqrt{\psi^{-1}(\lambda_{j})}-1\right)=f(\eta^{j+1}),$
where we recall that
$f(x)=f_{\psi}(x)=\psi(\lambda^{*}(x))\left(\sqrt{\psi^{-1}(\lambda^{*}(x))}-1\right).$
Then we can rearrange (B.26) to read
$\psi^{*}(\|S_{\tau}\|_{\widehat{V}_{\tau}^{-1}})\leq\psi^{*}(\eta^{j+1})\leq h_{\tau}(\lambda_{j})D_{\tau}(\delta_{j})+\eta^{j}\lambda_{j}(\eta-1)+f(\eta^{j+1}),$ (B.28)
where we’ve used that $\|S_{\tau}\|_{\widehat{V}_{\tau}^{-1}}\leq\eta^{j+1}$ on $E_{j}$ and the fact that $\psi^{*}$ is increasing. Now, we claim that
$\eta^{j}\lambda_{j}(\eta-1)\leq\frac{\psi^{*}(\eta^{j})}{2}.$
Let $s=\eta^{j}$ and define $R(s)=s\lambda_{j}(\eta-1)/\psi^{*}(s)$. We want to show that $R(s)\leq 1/2$. Using Lemma B.3, we have $\psi^{*}(s)\geq s\lambda^{*}(s)/2$. Moreover, if $\psi^{\prime\prime\prime}\geq 0$, then $\lambda^{*}(\eta s)\leq\eta\lambda^{*}(s)$, so
@@
$\psi^{*}(\|S_{\tau}\|_{\widehat{V}_{\tau}^{-1}})$ $\leq h_{\tau}(\lambda_{j})D_{\tau}(\delta_{j})+\frac{\psi^{*}(\eta^{j})}{2}+f(\eta^{j+1})$
$\leq h_{\tau}(\lambda_{j})D_{\tau}(\delta_{j})+\frac{\psi^{*}(\|S_{\tau}\|_{\widehat{V}_{\tau}^{-1}})}{2}+f(\eta\|S_{\tau}\|_{\widehat{V}_{\tau}^{-1}}),$
whence $\psi^{*}(\|S_{\tau}\|_{\widehat{V}_{\tau}^{-1}})\leq 2h_{\tau}(\lambda_{j})D_{\tau}(\delta_{j})+2f(2\|S_{\tau}\|_{\widehat{V}_{\tau}^{-1}})$, i.e.,
$\|S_{\tau}\|_{\widehat{V}_{\tau}^{-1}}\leq(\psi^{*})^{-1}\left(2h_{\tau}(\lambda_{j})D_{\tau}(\delta_{j})+2f(2\|S_{\tau}\|_{\widehat{V}_{\tau}^{-1}})\right).$ (B.29)
Now we return to the scenario when $\lambda_{j}=\psi^{-1}(1/\upsilon)$ in (B.27). In this case, we have
$\|S_{\tau}\|_{V_{\tau}^{-1}}\leq\frac{h_{\tau}(\psi^{-1}(1/\upsilon))D_{\tau}(\delta_{j})+g(\psi^{-1}(1/\upsilon))/\upsilon}{\psi^{-1}(1/\upsilon)}.$
@@
$B_{\tau}=\begin{cases}(\psi^{*})^{-1}\left(2h_{\tau}(\lambda^{*}(\eta^{j(\tau)+1})D_{\tau}(\delta_{j(\tau)})+2f(2\|S_{\tau}\|_{\widehat{V}_{\tau}^{-1}})\right),&\text{if }\lambda^{*}(\eta^{j(\tau)+1})\geq\psi^{-1}(1/\upsilon),\\ \frac{h_{\tau}(\psi^{-1}(1/\upsilon))D_{\tau}(\delta_{j})+g(\psi^{-1}(1/\upsilon))/\upsilon}{\psi^{-1}(1/\upsilon)},&\text{otherwise,}\end{cases}$ (B.30)
and $j(\tau)$ is the smallest $j\in\mathbb{N}_{0}$ such that $\|S_{\tau}\|_{V_{\tau}^{-1}}\leq\eta^{j+1}$. Therefore,
@@
Therefore, for large enough $t$, as long as $\psi^{-1}(1/\upsilon)<\lambda_{\max}$ (guaranteed by assumption), we have
@@
$\|S_{\tau}\|_{V_{\tau}^{-1}}\leq B_{\tau}\lesssim(\psi^{*})^{-1}\left(h_{\tau}(\lambda^{*}(\eta^{j(\tau)+1}))D_{\tau}(\delta_{j(\tau)})+f(2\|S_{\tau}\|_{V_{\tau}^{-1}})\right).$ (B.31)
It remains to bound $h_{\tau}(\lambda^{*}(\eta^{j(\tau)+1}))$. As was done in the proof of Proposition 4.4, notice that $\alpha_{t}\leq\gamma_{\max}(U_{0})/\gamma_{\min}(\widehat{V}_{t})$. Then,
$g_{\psi_{\tau}}(\lambda_{j})$ $=\sqrt{\psi^{-1}(\lambda_{j})}-\sqrt{\alpha_{t}\psi^{-1}(\lambda_{j})}$
$\geq\sqrt{\psi^{-1}(\lambda_{j}^{*})}\left(1-\sqrt{\frac{\gamma_{\max}(U_{0})}{\gamma_{\min}(\widehat{V}_{\tau})}}\right)$
$\geq\sqrt{\psi^{-1}(\psi^{-1}(1/\upsilon))}\left(1-\sqrt{\frac{\gamma_{\max}(U_{0})}{\gamma_{\min}(\widehat{V}_{\tau})}}\right).$